### v1

# The Value of Nothing111“Nowadays people know the price of everything and the value of nothing,” The Picture of Dorian Gray Wilde (1890): Multimodal Extraction of Human Values Expressed by TikTok Influencers

Abstract
Societal and personal values are transmitted to younger generations through interaction and exposure. Traditionally, children and adolescents learned values from parents, educators, or peers. Nowadays, social platforms serve as a significant channel through which youth (and adults) consume information, as the main medium of entertainment, and possibly the medium through which they learn different values. In this paper we extract implicit values from TikTok movies uploaded by online influencers targeting children and adolescents. We curated a dataset of hundreds of TikTok movies and annotated them according to the Schwartz Theory of Personal Values. We then experimented with an array of Masked and Large language model, exploring how values can be detected. Specifically, we considered two pipelines – direct extraction of values from video and a 2-step approach in which videos are first converted to elaborated scripts and then values are extracted.
Achieving state-of-the-art results, we find that the 2-step approach performs significantly better than the direct approach and that using a trainable Masked Language Model as a second step significantly outperforms a few-shot application of a number of Large Language Models. We further discuss the impact of fine-tuning and compare the performance of the different models on identification of values present or contradicted in the TikTok. Finally, we share the first values-annotated dataset of TikTok videos. Our results pave the way to further research on influence and value transmission in video-based social platforms.
1 Introduction
The transmission of values between generations is a key process for the preservation of society (Schönpflug 2001). Classic research assumed that transmission of values is carried out mainly by parents (Grusec 2011) or schools (Berson and Oreg 2016). However, youth spend hours every day on social media (Smahel et al. 2020), where they encounter new role models (Tolbert and Drogos 2019). The social media environment has been shown to strongly influence children’s behavior, in the context of consumer behavior (Coates et al. 2019). However, influencers in social media promote more than products: they form communities, by establishing connections, encouraging engagement, and promoting shared values (Himelboim and Golan 2023).
The Schwartz Theory of Personal Values (Schwartz 1992) offers a well-studied framework for the investigation of human motivation. It defines values as abstract goals, describing the end states aspired by people to achieve in their lives, such as success, independence, and care. These goals cover a wide range of motivations, with different individuals and societies, varying in the prioritization they place on different values (Schwartz 1992; Sagiv and Schwartz 2022). The theory identifies 19 core values, characterized by the goals they promote (Schwartz 2012). The values and their definitions, as used in the shared taskes at SemEval-2023 (Kiesel et al. 2022) and CLEF-2024 (Kiesel et al. 2024b) are provided in Table 1. Those 19 values are expressed through a wide variety of judgments, evaluations, and behaviors (Sagiv and Roccas 2021). Personal values, as defined by Schwartz, were consistently identified in different contexts, such as national cultures throughout the world (Sagiv and Schwartz 2022), organizational cultures (Daniel, Arieli, and Akerman 2024), traditional media (Kiesel et al. 2024a), and online disagreements (van der Meer et al. 2023).
In this work we suggest that social media influencers targeting youth can be characterized not only by genre, topic, and style, but also by the values they pursue and promote. For example, influencers hosting competitions between followers promote values of achievement, those presenting a unique craft project promote creativity, and others who pull pranks on strangers promote disregard for social norms. A specific example, presented in Table 2, depicts a gaming influencer failing to return tennis balls served by a professional tennis player. The values manifested in the video are Achievement (trying to properly return fast serves) and conflicted Face (the influencer show himself failing again and again to no avail).
How can we identify the values transmitted to youth through social media? This question has only recently began to be addressed. Some studies identified values prevalent in written social media using methods ranging from LIWC word matching to masked language models (e.g., BERT) (Kumar et al. 2018; Silva, Lo, and Lim 2020). However, recent trends lean heavily toward visual based social media, such as TikTok (Smahel et al. 2020; Rejeb et al. 2024). The shift toward visual mediums introduces further challenge, since it requires analyzing a clip, combining text, audio, and video. Further, the clips are typically very short, providing content through little stimuli. To the best of our knowledge, this is the first attempt to extract values from TikTok specifically, and visual social media in general.
In this work we experiment with an array of algorithms to extract values from TikTok videos targeting adolescents. To that end, we have curated a unique annotated corpus of 890 TikToks, posted by influencers targeting children and adolescents, covering a variety of genres prevalent in TikTok (see Section 3 for details). The corpus is useful for the investigation of values prevalent in this platform, and will be open for analysis.
We further experiment with and compare methods of value extraction, including fine-tuned Masked Language Models and Large Language Models. We go on to evaluate each method against the manual annotations, to identify the most effective one. Specifically, we consider two main pipelines: a direct extraction from LLMs, leveraging their multimodal capabilities, and a 2-step extraction approach in which a video is first converted to a detailed textual script (master scene heading, action, dialogue) and the values are extracted from the text, rather than directly from the video. We achieve a higher F-score than previously reported state-of-the-art results (Schroter, Dementieva, and Groh 2023; Legkas et al. 2024), in spite the more challenging setting (a direct comparison is unattainable, see Section 2). We find that that the 2-step approach perform better than the direct approach and that using a trainable Masked Language Model in the second step significantly outperforms a few-shot application of a Large Language Model. We further discuss the impact of fine-tuning and compare the performance of the different models on the present manifestation of values and the conflicted manifestations. Finally, we share the first values-annotated dataset of TikTok videos.
Value                                         Definition
SELF-DIRECTION–THOUGHT Freedom to cultivate one’s own ideas and abilities
SELF-DIRECTION–ACTION Freedom to determine one’s own actions
STIMULATION Excitement, novelty, and change
HEDONISM Pleasure and sensuous gratification
ACHIEVEMENT Success according to social standards
POWER–DOMINANCE Power through exercising control over people
POWER–RESOURCES Power through control of material and social resources
FACE Security and power through maintaining one’s public image and avoiding humiliation
SECURITY–PERSONAL Safety in one’s immediate environment
SECURITY–SOCIETAL Safety and stability in the wider society
TRADITION Maintaining and preserving cultural, family, or religious traditions
CONFORMITY–RULES Compliance with rules, laws, and formal obligations
CONFORMITY–INTERPERSONAL Avoidance of upsetting or harming other people
HUMILITY Recognizing one’s insignificance in the larger scheme of things
BENEVOLENCE–DEPENDABILITY Being a reliable and trustworthy member of the in-group
BENEVOLENCE–CARING Devotion to the welfare of in-group members
UNIVERSALISM–CONCERN Commitment to equality, justice, and protection for all people
UNIVERSALISM–NATURE Preservation of the natural environment
UNIVERSALISM–TOLERANCE Acceptance and understanding of those who are different from oneself
Table 1: Value definitions, Schwartz Theory of Personal Values (Schwartz 1992, 2012)
2 Related Work
TikTok and Influencers Research
Youth spend hours on various social media sites (Smahel et al. 2020; Tome and Cardita 2021), with TikTok gaining increased popularity (Anderson, Faverio, and Gottfried 2023; Woodward 2024). TikTok is characterized by engaging short videos that can include pervasive use of filters, and background music (Rejeb et al. 2024). The use of a costume “for you” page and “like” button increases immersion in the application (Montag, Yang, and Elhai 2021).
Social media in general, and TikTok specifically, are characterized by a large number of Social Media Influencers (SMIs). Social media influencers are social media users who gained large following (Ki et al. 2020) by creating and curating content that portrays an appealing online personality (Tafesse and Wood 2021). Influencers form parasocial relationships with their followers: members of the audience feels close, as if they know the influencers intimately (Reinikainen et al. 2020). Although parasocial relationships are known within traditional mass media, they may increase in the context of social media, because of its interactive nature (Conde and Casais 2023; Lou and Yuan 2019), and the disclosure of the influencers’ personal lives (Conde and Casais 2023). Followers often identify with their favorite influencers (Lou and Yuan 2019), making them likely role models for young viewers.
Ample research supports the understanding that children and adolescents can learn from mass traditional media (Prot et al. 2015), as well as social media (Ragelienė and Grønhøj 2021; Craig, Brooks, and Bichard 2021). From the media, youth can acquire information (Dorey and McCool 2009), but also form their worldviews (Duffett 2017). A recent study established that the following of social media influencers is associated with higher importance ascribed to materialistic values among Finish youth (Tuominen et al. 2023). We suggest that in order to further investigate the role of influencers in youth value transmission, we must identify the values they embrace.
Personal Values
Values are abstract goals describing end-states individuals aspire to achieve in their lives (Sagiv and Schwartz 2022; Schwartz 1992). Values direct a wide variety of behaviors, ranging between life-altering choices, such as career preferences (Lipshits-Braziler, Arieli, and Daniel 2024), and daily behaviors, such as the choice of touristic destinations (Saji, Vij, and Kabiraj 2015).
The Schwartz Theory of Personal Values is the leading theory of values in social psychology, well-validated in hundreds of samples across the world (Sagiv and Roccas 2017). Importantly, the theory extends beyond moral values, such as justice and care, to define a variety of motivations that can motivate individuals (Twito-Weingarten and Knafo‐Noam 2022), including, for example, safety, excitement, and influence. Moreover, the values are organized in a dynamic system of inter-relations, as the pursuit of one value may contrast the pursuit of conflicting values, yet enhance the pursuit of compatible ones (Schwartz 1992). These inherent trade-offs require individuals to prioritize among values. The inter-relations can be summarized into two main axes. One describes the opposition between self-transcendence (care for others, close and distant) and self-enhancement (focus on advancement of self-interest). The second describes the opposition between conservation (preservation of the status-quo) and openness to change (embracing new experiences).
Research established that children and adolescents hold meaningful values, following the content and structure defined by the Schwartz Theory of Personal Values (Daniel, Misgav, and Chomsky 2024; Daniel and Benish-Weisman 2019; Knafo-Noam, Daniel, and Benish-Weisman 2023). It is often agreed that children’s values are mainly learned from the social environment, in a process of value transmission (Twito and Knafo-Noam 2020). However, the social environment is varied and includes many socialization agents. Such agents are parents (Grusec 2011; Knafo-Noam, Barni, and Schwartz 2020), schools (Berson and Oreg 2016; Auer et al. 2023), the peer group (Benish-Weisman, Oreg, and Berson 2022; Benish‐Weisman 2024), and the national culture (Daniel et al. 2012). As social media becomes an important factor in the lives of adolescents, it is crucial to learn whether individuals within this sphere, such as influencers, become socialization agents teaching youth personal values.
Values in the Media
Schwartz values were identified in content displayed across social media platforms, such as Instagram (Trillò et al. 2021), Reddit (Chen et al. 2014), Facebook and Twitter (Kumar et al. 2018). Values in social media content were presented in varied ways, such as memes (Shifman 2019), and new-years’ resolutions (Hallinan et al. 2023). However, little is known about the prevalence of Schwartz values in youth-directed media. Some research exists on the prevalence of values in child-directed mass traditional media. An analysis of television shows addressing 9-11 year olds, showed high representation of values of fame (Uhls and Greenfield 2011). Similarly, television shows addressing young children, displayed work as an expression of a large variety of values, including benevolence, security, conformity, achievement, universalism, stimulation and hedonism (Aharoni, Te’eni-Harari, and Eyal 2024). Our work aims at opening the door to expand this work by developing an AI-based value classification in TikTok.
AI-Based Value Classification
The field of automated value detection has evolved significantly. Focusing on methods for classifying values according to Schwartz’s Theory of Personal Values (Schwartz 2012), early approaches leveraged the Linguistic Inquiry and Word Count (LIWC) (Pennebaker et al. 2015), or a targeted dictionaries (Ponizovskiy et al. 2020) detecting values through lexical mathcing. Values were detected in texts produced across various platforms, including Twitter (Kadić-Maglajlić, Lages, and Pantano 2024), narrative texts (Fischer et al. 2022), song lyrics (Demetriou et al. 2024), and social media platforms such as Reddit (Oulahbib, Benhaddi, and El Hadaj 2024) and Facebook (Kumar et al. 2018). Notable advances include the development of stack models combining LIWC features trained on Facebook data with cross-platform validation on Twitter (Silva, Lo, and Lim 2020). VALUENET (Qiu et al. 2022), introduced a large-scale text dataset and implemented a Transformer-based value regression model for human value modeling in dialogue systems.
Recent developments in predicting values were introduced with the application of masked langauge models. The SemEval-2023 (Kiesel et al. 2022) and CLEF-2024 (Kiesel et al. 2024b) competitions focused on value detection in argumentative texts and multi-lingual traditional media contexts respectively. These competitions demonstrated the effectiveness of BERT-based architectures (Schroter, Dementieva, and Groh 2023; Legkas et al. 2024) in capturing nuanced value expressions across two different text genres and languages. It also demonstrated that the algorithm developed in one context was not successful in identification of values in a second context (Kiesel et al. 2024b), asserting the importance of contextual models in value detection.
Generative AI methods were hardly utilized for value detection, with the exception of the FULCRA project (Yao et al. 2024), leveraging LLM-generated text pairs for safety-oriented value detection. Our work extends these methodological advances by pioneering value detection in TikTok’s unique multimodal environment. Unlike previous approaches that focused on single modalities, TikTok content presents a complex interplay of written text, audio, and visual elements. To address this challenge, we conduct a comparative analysis of two leading approaches: BERT-based architectures and LLM-based methods. By evaluating these distinct methodologies on the same TikTok dataset, we provide the first systematic comparison of their effectiveness in detecting values within short-form video content. This novel application to multimodal social media content, combined with our comparative analysis, can contribute valuable insights to both social media analysis and automated value detection research, particularly for platforms where meaning is conveyed through multiple channels simultaneously.
Figure 1: Values per movie in the annotated data.
Figure 2: Occurrences of each value in the annotated dataset. Blue bars indicate the appearance of a value (e.g., achievement) and red bars indicate the conflicted value (e.g., apathy and lack of ambition, conflicting with achievement).
3 Data and Annotation
3.1 Data Curation
We identified potential TikTok influencers using surveys of children and adolescents (anonymous survey, approved by a the local IRB), and lists provided by three large language models (ChatGPT, Gemini, Bing). Two research assistants reviewed the potential influencers, to confirm they are targeting children or adolescents and to classify them by genre. This process resulted in a list of 89 influencers, producing content in the major genres prevalent in child- and adolescent-targeted TikTok: beauty and skincare, lifestyle, entertainment and pranks, crafts and DYI, vlogging, gaming and dancing, singing and lip-syncing. We sampled 10 TikToks from the page of each influencer. We first identified the pinned TikToks, and supplemented the sample with the last TikToks available on March 19, 2024, until a total of 10 TikToks was reached from each influencer, to the total N = 890.
We only selected TikToks that include verbal sound, such as talking, narration or song. The crawling and curation were done with slight modifications to the API shared by (Teather 2024).
3.2 Annotation Protocol
We constructed a codebook on the basis of the Schwartz theory of personal values (Schwartz 1992, 2012). The codebook included a definition to each of the 19 values derived from the theory of personal values (Schwartz 2012). In addition, the theory included examples of expression of these values that are common on TikTok. The codebook included instructions for identifying values as present (1, marked as +), when it is expressed or pursued in the TikTok; as absent (0), when the TikTok does not refer to the value; or as conflicted (-1, marked as -), when the TikTok contradicts or hinders the pursuit of the value. Each TikTok could be annotated as referring to as many or as few values as relevant. A team of annotators was trained in value annotation. Each TikTok was independently annotated by two annotators. A third annotator consolidated the annotations by resolving discrepancies. To assess inter-rater agreement on our nominally scaled data, we considered various metrics beyond traditional $\pi$ and $\kappa$ statistics, which are known to be sensitive to imbalanced datasets (Gwet 2008). Given that our dataset exhibits significant imbalance, with the majority category (not present) accounting for 0.76 of annotations and rare categories (conflict and dominant) comprising only 0.06 and 0.03, we calculated Gwet’s AC1 coefficient, which has been demonstrated to be more robust against prevalence effects and marginal probability issues than Cohen’s Kappa (Wongpakaran et al. 2013). Our analysis yielded an AC1 score of 0.77, indicating substantial inter-rater agreement despite the challenging data distribution. The codebook is available at Appendix A.1.
In total, the annotated dataset of 890 TikTok TikToks, were annotated with a total of 2968 labels (of values and conflicting values). The distribution of values per movie is presented in Figure 1 and the frequency of each value in our dataset is presented in Figure 2. We make the annotated dataset available at [URL to added in camera ready version].
4 Computational Approach
Our computational approach could be placed on three axes: (i) Masked vs. Auto-regressive Language Models, (ii) Direct vs. indirect, 2-step extraction from TikTok videos (marked d versus s, repestecively), and (iii) Using pretrained LLMs vs. fine-tuned models (marked FT). In the remainder of this section we describe each axis and the way we integrate them. Figure 3 provides a high-level illustration of the architectures.
Figure 3: Direct and indirect extraction of values from TikTok movies. Fine tuning the textual models on scripts is optional. LLMs are used in a zero/few-shot manner. The MLM must be trained on a subsample in a supervised manner.
Masked vs. Auto-regressive Language Models
The classic machine learning framework depends on the availability of a labeled dataset, facilitating training in a supervised manner. A Masked Language Model (MLM) is pre-trained to predict words that are masked at random. The MLM is then used to create a contextual embedding of an input text, pushing it forward into a classification head. In this work we experiment with both roBERTa-large (Liu 2019) and DeBERTaV2-xlarge (He, Gao, and Chen 2023) with an additional classification header performing transfer learning (Devlin 2018). We use this supervised approach on textual data only – the scripts and director notes extracted from the videos by another model.
Generative (auto-regressive) Large Language Models (LLMs) have been demonstrated to perform well on a range of language processing and multimodal tasks. LLMs do not require specific training (transfer learning) as classification is done by prompting the model with the task description (Radford et al. 2019; Brown et al. 2020; Kojima et al. 2022). State-of-the-Art models such as ChatGPT.4o (OpenAI 2024) and Gemini 1.5 Pro (Google 2024) are multimodal - allowing the user to process different modalities such as text, sound and video. In this work we use Gemini 1.5 Pro with a prompt tailored for the specific tasks of script extraction and for direct and 2-step classification (see below). In the 2-step setting we also use Mistral 7B (Jiang et al. 2023).
Specifically, the prompt contains a short formal description of Schwartz’s theory of basic values, a definition for each value, along with some examples and a request to identify the values in a given TikTok (an .mp4 file is uploaded). The exact prompts are available in Appendix A.
Direct vs. 2-step Extraction
We use an LLM as a classifier in two settings – direct and indirect (2-steps). In the direct setting we prompt the model to identify and extract the values directly from the TikTok video. In the indirect approach we first prompt an LLM to extract the full script and director’s notes from a TikTok and then extract the values from the script, either in a supervised manner using one of the MLMs or in a few-shot manner, using the exact same prompt we used in the direct approach, albeit, replacing the word ‘movie’ by the word ‘script’.
An example of a generated script is presented in Table 2.
We experiment with the different approaches for three main reasons: First, previous work reported results over text in a supervised setting. Secondly, the inner-workings of (multimodal) LLMs are opaque and different prompts and tasks may result in unstable results. Finally, while LLMs are flashy and do not require supervision, it was argued that they underperform on nuanced tasks such as the detection of emotions, figurative language, implicit hate etc. (Ziems et al. 2024).
Genre: Sports/Challenge
Sound: Yes
NARRATOR: "This pro tennis player Taylor Fritz with one of the fastest serves on tour. His fastest serve ever was 240 KM/H, but can I, a person who’s never played tennis, return one?"
[pro]
TAYLOR FRITZ is shown on a tennis court serving, then staring into the camera.
[#9 IN THE WORLD]
[with one]
[fastest]
[serves]
[240KM/H]
[I]
[never]
[return]
The NARRATOR is shown holding a tennis racket, about to return a serve.
[SERVE 1]
TAYLOR FRITZ serves the ball. The NARRATOR misses.
NARRATOR: "Oof"
[SERVE 4]
TAYLOR FRITZ serves. The NARRATOR misses again.
NARRATOR: "Oh, yo, that was..."
[SERVE 10]
TAYLOR FRITZ serves. The NARRATOR misses.
NARRATOR: "F*ck"
[SERVE 18]
TAYLOR FRITZ serves. The NARRATOR misses.
NARRATOR: "F*ck. [giggles] Yo!"
[SERVE 24]
TAYLOR FRITZ serves. The NARRATOR misses.
NARRATOR: "F*ck, I thought I could do this."
[SERVE 32]
TAYLOR FRITZ serves. The NARRATOR misses.
NARRATOR: "F*ck no, I cannot!"
[SERVE 38]
TAYLOR FRITZ serves. The NARRATOR misses.
[SERVE 47]
TAYLOR FRITZ serves. The NARRATOR hits the ball back.
NARRATOR: "Oh! Where’d it go?"
[50TH SERVE]
TAYLOR FRITZ serves for the last time.
NARRATOR: "Oh, come on."
Table 2: The script generated by the video-to-script prompt (Appendix A.2) for video ID 7341895431883967746
Transfer learning and Fine-tuning
Pretrained Language Models, either masked or auto-regressive, serve as foundational models and can be used as-is, either by prompting (LLMs) or by performing transfer learning on a small annotated sample (MLMs). However, both types of language models could be further fine-tuned on in-domain data, before the classification phase. Typically [^2] Fine-tuning can also be done as part of the classifier training., fine-tuning is done in a similar way to the pretraining, that is – given a dedicated corpora, masked language models are fine-tuned by predicting randomly masked words (Zaken, Goldberg, and Ravfogel 2022), and autoregressive models are fine-tuned on next-word prediction task (Hu et al. 2021). In this work we use the roberta-large framework to fine-tune the roBERTa-large, the DeBERTaV2-xlarge, and the Mistral LLM. We report results obtained both by the basic pretrained models and the fine-tuned models.
Compute Resources
All experiments and training were executed on Google Colab Enterprise runtime environment with an NVIDIA Tesla A100 GPU (a2-highgpu-1g configuration).
5 Results, Analysis, and Discussion
The F-scores achieved by each of the models on each of the nineteen values are presented in Figure 4 and the weighted and macro F-scores are presented in Table 3. The best overall results for values and their conflicted counterparts were achieved with the non-fine-tuned RoBERTa-large (roBERTas) with weighted F-scores of 0.49/0.42/0.47 (present/contradicted/all [$+/-/all$], respectively). These levels of classification reflect the opacity of the signal, that corresponds to a multidimensional and complex human phenomenon. In fact, even human annotators required extensive training and still found the task non-trivial. At the same time, these results are competitive with previously reported state-of-the-art results (Schroter, Dementieva, and Groh 2023; Legkas et al. 2024) achieved in less challenging settings in which (i) text, rather than video, was the original medium of the input, and (ii) the number of examples available for training was an order of magnitude larger. [^3] A direct comparison is unattainable due to the very different nature of the data in terms of source, medium and style.
Looking at specific values (Figure 4(a)), best classification was achieved for Achievement (0.71), Benevolence Dependability (0.61), and Security Personal (0.58), all by the RoBERTas model. Achievement, present in more than one third of the videos, is the most frequent value in the annotated dataset, thus more occurrences are found in training and in the test set. However, the two other values occur far less than most values, hence performance cannot be attributed to prior frequency alone.
A fine-grained analysis of the results by method and value reveals interesting patterns, suggesting a number of insights. We address some of these patterns in the remainder of this section.
(a) Present manifestations (Max. F-Score: 0.71)
(b) Contradicting manifestations (Max. F-Score: 0.54)
Figure 4: F-Scores per each model and value. Values appearing less than 30 times were not included in the BERT based MLM models, thus excluded from the analysis (appearing with score 0 in the figure).
Masked Language Models Large Language Models
[…]
Weighted+ 0.49 0.44 0.30 0.33 0.29 0.33 0.40
Macro+ 0.44 0.38 0.26 0.29 0.33 0.28 0.33
Weighted- 0.42 0.41 0.33 0.30 0.17 0.21 0.22
Macro- 0.35 0.35 0.27 0.25 0.22 0.18 0.22
Weightedall 0.47 0.43 0.31 0.33 0.26 0.30 0.35
Macroall 0.39 0.36 0.25 0.26 0.29 0.24 0.29
Table 3: Weighted and Macro F-Score results for the values ($+$), the conflicted values (--) and all ($all$). $FT$: fined-tuned model; $s$ indicates extraction from scripts (2-step extraction); $d$ indicates direct extraction from video.
MLMs vs. LLMs
The best LLM results were obtained by Geminid. However, it was outperformed by both RoBERTas and DeBERTas in all metrics and over the three partitions ($\{weighted,macro\}\times\{+,-,all\}$, Table 3). Looking at single values, we observe that Gemini outperforms the MLMs in cases in which the number of occurrences of the value is minimal, hence supervised training is hampered, see results per value in Figure 4 and compare to the value distribution Figure 2. An example is Universalism Concern in the present manifestations, and Power Resources in the contradiction manifestation.
These results are consistent with the literature, reporting that strong supervised models tend to outperform current LLMs on tasks requiring nuance, typical to the Computational Social Science (CSS) research (Ziems et al. 2024).
Pretraining vs. Fine-tuning
All the fine-tuned models fell short, comparing to their pretrained counterparts. While somewhat surprising, previous studies report similar trends in various tasks and settings, e.g., (Zaken, Goldberg, and Ravfogel 2022; Kumar et al. 2022), as well as in the best performer in she value detection shared task (Schroter, Dementieva, and Groh 2023). Given the task and data at hand, we attribute this result to the combination of two main factors: (i) The relatively small dataset used for finetuning, (ii) The large number of imbalanced classes, as presented in Figure 2. Future work will explore the impact of each of these factors.
Direct vs. 2-step Extraction
The textual models (MLMs) applied in a 2-step pipeline and trained in a supervised manner, achieved better results than Geminid applied directly on the videos. However, comparing apples to apples, considering only the LLMs, we found that the direct application of Geminid on the videos significantly outperformed the application of both Geminis and Mistrals on the scripts in the 2-step pipeline. These results highlight the importance of the visual and sound elements to the TikTok platform, and thus the potential of the direct multimodal approach.
Present vs. Conflicting Manifestations
Most present values are more frequent (Figure 2) and better detected than values in conflict (Figure 4 and Table 3). It is interesting to note that while Geminid, our top performing LLM, showed decent performance on the present values (Weighted+ F-score of 0.4, compared to 0.49 by the top performing model and 0.33 by the second best LLM), it perform poorly on the conflicting values, achieving an F-score of only 0.22, compared to the 0.42 by the top performing model and only slightly better than the other LLMs. This result is surprising: while MLM results are expected to drop due to the smaller number of labeled instances available for training, we would have expected LLMs, used in a few-shot manner, to perform comparably on present and conflicting values. We have no plausible explanation for this significant drop in performance and intend to further explore it in the future.
Video to Script Conversion
Finally, the use of the multimodal Gemini 1.5Pro to convert videos to detailed textual script (master scene heading, action, dialogue) proved beneficial for the task. We find the conversion to be reasonably accurate and the level of details provide useful markers. This conversion was vital for the application of the successful MLM. Two points are worth highlighting: First, before converging on the final prompt (see Appendix A.2) we have experimented with different prompts, producing different outputs in various formats and detail. The evaluation of the quality of the script was done in a qualitative manner by two of the authors. Second, even the state-of-the-art models often fail to capture important details. For example, in a humerus nonsensical TikTok showing the influencer air-driving a car as a part of a guessing charade game, practically air driving, floating around in the room and bumping into objects, the generated script failed to reflect the fact that the influencer was actually moving in the air – the main point of the video.
Limitations
This study is not without limitations. First, we sampled influencers in seven specific genres. In our dataset, we found much variability between the genres in content as well as style. We identified genres most important for children and adolescents, yet know that TikTok is extremely varied, and other genres exist. Thus, it is yet to verify that our results generalize across TikTok. Second, we applied the Schwartz personal values theory to define and identify values. This theory was tested across cultures and in hundreds of samples across the world (Sagiv and Schwartz 2022), and is considered the leading theoretical framework of values in social science. At the same time, the theory was not constructed specifically within the context of social media, and does not focus on unique aspects of this context. Specialized theories can add information for future investigation (Shifman et al. 2025). Last, in our analysis, we utilized two MLMs, and two LLMs. the use of other models may offer different results, and future studies can extend this investigation further.
6 Conclusions and Implications
We present pipeline for the extraction of personal values from TikTok, and demonstrate their utility. We show that (supervised) BERT based models outperform LLMs in value extraction, and offer a method for conversion of multimodal TikToks into text appropriate for BERT training and analysis. We also show variability in detection between different values, and between values supported and those contradicted in the TikToks.
The methods presented have important implications for research and society. TikTok emerged as a leading platform among youth (Anderson, Faverio, and Gottfried 2023; Woodward 2024), yet at the same time produces worry among parents, educators, and decision makers (Conte et al. 2024). By applying value detection, research can move on to investigate value transmission in TikTok at scale. Such research can provide society with much needed information regarding the social environment of young people. It is important to note that advertisers and influencers can make similar use of the method, in order to target youth more effectively. We follow (De Veirman, Hudders, and Nelson 2019) in call for policy makers and educators to mitigate misuse of online platforms by enabling youth to better identify patterns of influence in social media.
Appendix A Prompts and Responses
In this appendix we provide the prompts used for value extraction (direct and 2-step extraction) and the prompt used in converting a movie to textual script, along with an example of a generated script and a the values extracted from it.
A.1 Prompt for Value Extraction
The prompt used for direct extraction is provided at Figures 5 and 6. This very same “prompt” served as the annotation guidelines (code book) offered to annotators. Note that the figure contains an examples (“few shot”) for only one value (SELF-DIRECTION-THOUGHT) while the actual prompt (and the annotation guidelines) provided examples for each of the values. The same prompt was used for few-shot extraction from scripts in the 2-step extraction, the sole difference is that the opening line is “You are given a movie script of a TikTok video”, rather than “You are given a TikTok video”. The full code book (prompt) is available at [add URL].
Figure 5: Annotation guidelines and prompt for direct extraction.
Figure 6: Additional guidelines, appended after all the examples (code book).
A.2 Prompt for Srtipt Extraction
The prompt used to convert TikTok movies to textual scripts is available in Figure 7; the script generated by applying this prompt on TikTok ID 7341895431883967746 is presented in Table 2. The manual annotations for this specific movie are: Achievement- present (1), Face - conflict (-1).
Figure 7: The prompt used to convert TikTok movies to textual scripts.

### v2

@@
-The Value of Nothing111“Nowadays people know the price of everything and the value of nothing,” The Picture of Dorian Gray Wilde (1890): Multimodal Extraction of Human Values Expressed by TikTok Influencers
+The Value of Nothing111“Nowadays people know the price of everything and the value of nothing,” The Picture of Dorian Gray (Wilde 1890): Multimodal Extraction of Human Values Expressed by TikTok Influencers
 Abstract
@@
 In this paper we extract implicit values from TikTok movies uploaded by online influencers targeting children and adolescents.
-We curated a dataset of hundreds of TikTok movies and annotated them according to the Schwartz Theory of Personal Values.
-We then experimented with an array of Masked and Large language model, exploring how values can be detected.
-Specifically, we considered two pipelines – direct extraction of values from video and a 2-step approach in which videos are first converted to elaborated scripts and then values are extracted.
-Achieving state-of-the-art results, we find that the 2-step approach performs significantly better than the direct approach and that using a trainable Masked Language Model as a second step significantly outperforms a few-shot application of a number of Large Language Models.
-We further discuss the impact of fine-tuning and compare the performance of the different models on identification of values present or contradicted in the TikTok.
+We curated a dataset of hundreds of TikTok movies and annotated them according to the well established Schwartz Theory of Personal Values.
+We then experimented with an array of language models, investigating their utility in value identification.
+Specifically, we considered two pipelines: direct extraction of values from video and a 2-step approach in which videos are first converted to elaborated scripts and values are extracted from the textual scripts.
+We find that the 2-step approach performs significantly better than the direct approach and that using a few-shot application of a Large Language Model in both stages outperformed the use of a fine-tuned Masked Language Model in the second stage.
+We further discuss the impact of continuous pretraining and fine-tuning and compare the performance of the different models on identification of values endorsed or confronted in the TikTok.
 Finally, we share the first values-annotated dataset of TikTok videos.
-Our results pave the way to further research on influence and value transmission in video-based social platforms.
+To the best of our knowledge, this is the first attempt to extract values from TikTok specifically, and visual social media in general.
+Our results pave the way to future research on value transmission in video-based social platforms.
 1 Introduction
@@
 However, youth spend hours every day on social media (Smahel et al.
-2020), where they encounter new role models (Tolbert and Drogos 2019).
+2020), where they find new role models (Tolbert and Drogos 2019).
 The social media environment has been shown to strongly influence children’s behavior, in the context of consumer behavior (Coates et al.
 2019).
-However, influencers in social media promote more than products: they form communities, by establishing connections, encouraging engagement, and promoting shared values (Himelboim and Golan 2023).
+However, influencers on social media promote more than products: they form communities and promote shared values (Himelboim and Golan 2023).
 The Schwartz Theory of Personal Values (Schwartz 1992) offers a well-studied framework for the investigation of human motivation.
-It defines values as abstract goals, describing the end states aspired by people to achieve in their lives, such as success, independence, and care.
-These goals cover a wide range of motivations, with different individuals and societies, varying in the prioritization they place on different values (Schwartz 1992; Sagiv and Schwartz 2022).
-The theory identifies 19 core values, characterized by the goals they promote (Schwartz 2012).
-The values and their definitions, as used in the shared taskes at SemEval-2023 (Kiesel et al.
-2022) and CLEF-2024 (Kiesel et al.
-2024b) are provided in Table 1.
-Those 19 values are expressed through a wide variety of judgments, evaluations, and behaviors (Sagiv and Roccas 2021).
-Personal values, as defined by Schwartz, were consistently identified in different contexts, such as national cultures throughout the world (Sagiv and Schwartz 2022), organizational cultures (Daniel, Arieli, and Akerman 2024), traditional media (Kiesel et al.
-2024a), and online disagreements (van der Meer et al.
+It defines values as abstract goals, describing the end states people aspire to achieve in their lives, such as success, independence, and care.
+These goals cover a wide range of motivations, with different individuals and societies varying in the prioritization they place on different values (Schwartz 1992; Sagiv and Schwartz 2022).
+Values are expressed through a wide variety of judgments, evaluations, and behaviors (Sagiv and Roccas 2021).
+The theory identifies 19 core values, characterized by the boader goals they convey or serve (Schwartz 2012).
+Personal values, as defined by Schwartz (see Table 1), were consistently identified in different contexts, such as national cultures throughout the world (Sagiv and Schwartz 2022), organizational cultures (Daniel et al.
+2024a), traditional media (Kiesel et al.
+2024a), and online discourse (van der Meer et al.
 2023).
-In this work we suggest that social media influencers targeting youth can be characterized not only by genre, topic, and style, but also by the values they pursue and promote.
-For example, influencers hosting competitions between followers promote values of achievement, those presenting a unique craft project promote creativity, and others who pull pranks on strangers promote disregard for social norms.
-A specific example, presented in Table 2, depicts a gaming influencer failing to return tennis balls served by a professional tennis player.
-The values manifested in the video are Achievement (trying to properly return fast serves) and conflicted Face (the influencer show himself failing again and again to no avail).
+In this work, we suggest that social media influencers targeting youth, can be characterized not only by genre, topic, and style, but also by the values they endorse and the values they confront and disregard.
+To list a number of examples, an influencer hosting a competition between followers endorses values of achievement; an influencer presenting a unique craft project endorses values of creativity; influencers who pull pranks on strangers confront the values of respect for social norms.
+A number of values can be endorsed/confronted in the same piece of content.
+For example, in one video (script available in Table 2), a gaming influencer fails to return tennis balls served by a professional tennis player.
+The value of Achievement (trying to properly return fast serves) is endored in the video.
+At the same time, the value of Face is confronted, as the influencer is humiliated (shows himself failing again and again to no avail).
 How can we identify the values transmitted to youth through social media?
-This question has only recently began to be addressed.
-Some studies identified values prevalent in written social media using methods ranging from LIWC word matching to masked language models (e.g., BERT) (Kumar et al.
-2018; Silva, Lo, and Lim 2020).
-However, recent trends lean heavily toward visual based social media, such as TikTok (Smahel et al.
+Some recent studies identified values prevalent in written social media using a variety of methods e.g.
+(Kumar et al.
+2018; Silva et al.
+2020) ranging from linguistic inquiry and word count (LIWC), word matching to masked language models, such as BERT.
+However, current trends in social media lean heavily toward visual based media, such as TikTok (Smahel et al.
 2020; Rejeb et al.
 2024).
-The shift toward visual mediums introduces further challenge, since it requires analyzing a clip, combining text, audio, and video.
-Further, the clips are typically very short, providing content through little stimuli.
+The shift toward visual mediums introduces a further challenge, since it requires analyzing a clip in a multimodal way, combining text, audio, and video.
+In addition, the clips are typically very short, providing rich content through limited stimuli.
 To the best of our knowledge, this is the first attempt to extract values from TikTok specifically, and visual social media in general.
-In this work we experiment with an array of algorithms to extract values from TikTok videos targeting adolescents.
-To that end, we have curated a unique annotated corpus of 890 TikToks, posted by influencers targeting children and adolescents, covering a variety of genres prevalent in TikTok (see Section 3 for details).
+In this work we have curated and annotated a unique corpus of 885 TikToks, posted by influencers targeting children and adolescents, covering a variety of genres prevalent in TikTok (see Section 3 for details).
 The corpus is useful for the investigation of values prevalent in this platform, and will be open for analysis.
-We further experiment with and compare methods of value extraction, including fine-tuned Masked Language Models and Large Language Models.
-We go on to evaluate each method against the manual annotations, to identify the most effective one.
+We experiment with and compare an array of algorithms of value classification, including fine-tuned Masked Language Models and Large Language Models.
 Specifically, we consider two main pipelines: a direct extraction from LLMs, leveraging their multimodal capabilities, and a 2-step extraction approach in which a video is first converted to a detailed textual script (master scene heading, action, dialogue) and the values are extracted from the text, rather than directly from the video.
-We achieve a higher F-score than previously reported state-of-the-art results (Schroter, Dementieva, and Groh 2023; Legkas et al.
-2024), in spite the more challenging setting (a direct comparison is unattainable, see Section 2).
-We find that that the 2-step approach perform better than the direct approach and that using a trainable Masked Language Model in the second step significantly outperforms a few-shot application of a Large Language Model.
-We further discuss the impact of fine-tuning and compare the performance of the different models on the present manifestation of values and the conflicted manifestations.
-Finally, we share the first values-annotated dataset of TikTok videos.
+In spite of the multimodal complexity, we achieve a high F-score, higher than the state-of-the-art results reported on the task of extracting values from texts (Schroter et al.
+2023; Legkas et al.
+2024).
+We find that the 2-step approach performs better than the direct approach and that using a few-shot application of a Large Language Model outperforms a trainable Masked Language Model in the second step.
+We further discuss the impact of fine-tuning and compare the performance of the different models on the endorsed and the confronted values.
 Value                                         Definition
@@
 2 Related Work
-TikTok and Influencers Research
+The TikTok Influencers Ecosystem
 Youth spend hours on various social media sites (Smahel et al.
-2020; Tome and Cardita 2021), with TikTok gaining increased popularity (Anderson, Faverio, and Gottfried 2023; Woodward 2024).
-TikTok is characterized by engaging short videos that can include pervasive use of filters, and background music (Rejeb et al.
+2020), with TikTok gaining increased popularity (Anderson et al.
+2023; Woodward 2024).
+TikTok is characterized by engaging short videos that can include widespread use of filters, and background music (Rejeb et al.
 2024).
-The use of a costume “for you” page and “like” button increases immersion in the application (Montag, Yang, and Elhai 2021).
+The use of a custom “For You” page and “Like” button increases immersion in the application (Montag et al.
+2021).
 Social media in general, and TikTok specifically, are characterized by a large number of Social Media Influencers (SMIs).
-Social media influencers are social media users who gained large following (Ki et al.
+Social media influencers are social media users who gained a large following (Ki et al.
 2020) by creating and curating content that portrays an appealing online personality (Tafesse and Wood 2021).
-Influencers form parasocial relationships with their followers: members of the audience feels close, as if they know the influencers intimately (Reinikainen et al.
+Influencers form parasocial relationships with their followers: members of the audience feel close to influencers, as if they know them intimately (Reinikainen et al.
 2020).
-Although parasocial relationships are known within traditional mass media, they may increase in the context of social media, because of its interactive nature (Conde and Casais 2023; Lou and Yuan 2019), and the disclosure of the influencers’ personal lives (Conde and Casais 2023).
+Influencers are especially likely to form such relationships because of social media’s interactive nature (Conde and Casais 2023; Lou and Yuan 2019), and the disclosure of the influencers’ personal lives (Conde and Casais 2023).
 Followers often identify with their favorite influencers (Lou and Yuan 2019), making them likely role models for young viewers.
 Ample research supports the understanding that children and adolescents can learn from mass traditional media (Prot et al.
-2015), as well as social media (Ragelienė and Grønhøj 2021; Craig, Brooks, and Bichard 2021).
+2015), as well as social media (Craig et al.
+2021).
 From the media, youth can acquire information (Dorey and McCool 2009), but also form their worldviews (Duffett 2017).
@@
 Values are abstract goals describing end-states individuals aspire to achieve in their lives (Sagiv and Schwartz 2022; Schwartz 1992).
-Values direct a wide variety of behaviors, ranging between life-altering choices, such as career preferences (Lipshits-Braziler, Arieli, and Daniel 2024), and daily behaviors, such as the choice of touristic destinations (Saji, Vij, and Kabiraj 2015).
+Values direct a wide variety of behaviors, ranging between life-altering choices such as career preferences (Lipshits-Braziler et al.
+2024), and daily behaviors, such as the choice of touristic destinations (Saji et al.
+2015).
 The Schwartz Theory of Personal Values is the leading theory of values in social psychology, well-validated in hundreds of samples across the world (Sagiv and Roccas 2017).
-Importantly, the theory extends beyond moral values, such as justice and care, to define a variety of motivations that can motivate individuals (Twito-Weingarten and Knafo‐Noam 2022), including, for example, safety, excitement, and influence.
-Moreover, the values are organized in a dynamic system of inter-relations, as the pursuit of one value may contrast the pursuit of conflicting values, yet enhance the pursuit of compatible ones (Schwartz 1992).
+Importantly, the theory extends beyond moral values, such as justice and care, to define a variety of motivations that can drive individuals (Twito-Weingarten and Knafo‐Noam 2022), including, for example, safety, excitement, and influence.
+Moreover, the values are organized in a dynamic system of inter-relations, as the pursuit of one value may contrast the pursuit of conflicting values, and enhance the pursuit of compatible ones (Schwartz 1992).
 These inherent trade-offs require individuals to prioritize among values.
 The inter-relations can be summarized into two main axes.
-One describes the opposition between self-transcendence (care for others, close and distant) and self-enhancement (focus on advancement of self-interest).
-The second describes the opposition between conservation (preservation of the status-quo) and openness to change (embracing new experiences).
-Research established that children and adolescents hold meaningful values, following the content and structure defined by the Schwartz Theory of Personal Values (Daniel, Misgav, and Chomsky 2024; Daniel and Benish-Weisman 2019; Knafo-Noam, Daniel, and Benish-Weisman 2023).
+One describes the tension between self-transcendence (care for others, close and distant) and self-enhancement (focus on advancement of self-interest).
+The second describes the tension between conservation (preservation of the status-quo) and openness to change (embracing new experiences).
+The list of Schwartz values and their short description is available at Table 1.
+Research established that children and adolescents hold meaningful values, following the content and structure defined by the Schwartz Theory of Personal Values (Daniel et al.
+2024b; Daniel and Benish-Weisman 2019; Knafo-Noam et al.
+2023).
 It is often agreed that children’s values are mainly learned from the social environment, in a process of value transmission (Twito and Knafo-Noam 2020).
 However, the social environment is varied and includes many socialization agents.
-Such agents are parents (Grusec 2011; Knafo-Noam, Barni, and Schwartz 2020), schools (Berson and Oreg 2016; Auer et al.
-2023), the peer group (Benish-Weisman, Oreg, and Berson 2022; Benish‐Weisman 2024), and the national culture (Daniel et al.
+Such agents are parents (Grusec 2011; Knafo-Noam et al.
+2020), schools (Berson and Oreg 2016), the peer group (Benish-Weisman et al.
+2022; Benish‐Weisman 2024), and the national culture (Daniel et al.
 2012).
-As social media becomes an important factor in the lives of adolescents, it is crucial to learn whether individuals within this sphere, such as influencers, become socialization agents teaching youth personal values.
-Values in the Media
+As social media becomes an important factor in the lives of adolescents, it is crucial to learn whether individuals within this sphere, such as influencers, become socialization agents teaching personal values to youth.
+Values Social Media
 Schwartz values were identified in content displayed across social media platforms, such as Instagram (Trillò et al.
@@
 2018).
-Values in social media content were presented in varied ways, such as memes (Shifman 2019), and new-years’ resolutions (Hallinan et al.
+Values were presented in social media content in varied ways, such as memes (Shifman 2019), and new-years’ resolutions (Hallinan et al.
 2023).
@@
 Some research exists on the prevalence of values in child-directed mass traditional media.
-An analysis of television shows addressing 9-11 year olds, showed high representation of values of fame (Uhls and Greenfield 2011).
-Similarly, television shows addressing young children, displayed work as an expression of a large variety of values, including benevolence, security, conformity, achievement, universalism, stimulation and hedonism (Aharoni, Te’eni-Harari, and Eyal 2024).
-Our work aims at opening the door to expand this work by developing an AI-based value classification in TikTok.
+An analysis of television-shows addressing 9-11 year olds, showed high representation of fame values (Uhls and Greenfield 2011).
+Similarly, television shows addressing young children, displayed work as an expression of a large variety of values, including benevolence, security, conformity, achievement, universalism, stimulation and hedonism (Aharoni et al.
+2024).
+Our work aims at opening the door to extract values at scale by developing an AI-based value detection in TikTok movies.
 AI-Based Value Classification
@@
 Focusing on methods for classifying values according to Schwartz’s Theory of Personal Values (Schwartz 2012), early approaches leveraged the Linguistic Inquiry and Word Count (LIWC) (Pennebaker et al.
-2015), or a targeted dictionaries (Ponizovskiy et al.
-2020) detecting values through lexical mathcing.
-Values were detected in texts produced across various platforms, including Twitter (Kadić-Maglajlić, Lages, and Pantano 2024), narrative texts (Fischer et al.
+2015), or a targeted dictionary (Ponizovskiy et al.
+2020) detecting values through lexical matching.
+Values were detected in texts produced across various platforms, including Twitter (Kadić-Maglajlić et al.
+2024), narrative texts (Fischer et al.
 2022), song lyrics (Demetriou et al.
-2024), and social media platforms such as Reddit (Oulahbib, Benhaddi, and El Hadaj 2024) and Facebook (Kumar et al.
+2024), and social media platforms such as Reddit (Oulahbib et al.
+2024) and Facebook (Kumar et al.
 2018).
-Notable advances include the development of stack models combining LIWC features trained on Facebook data with cross-platform validation on Twitter (Silva, Lo, and Lim 2020).
+Notable advances include the development of stack models combining LIWC features trained on Facebook data with cross-platform validation on Twitter (Silva et al.
+2020).
 VALUENET (Qiu et al.
 2022), introduced a large-scale text dataset and implemented a Transformer-based value regression model for human value modeling in dialogue systems.
-Recent developments in predicting values were introduced with the application of masked langauge models.
-The SemEval-2023 (Kiesel et al.
+Recent developments in value prediction were introduced with the application of masked language models.
+The SemEval-2022 (Kiesel et al.
 2022) and CLEF-2024 (Kiesel et al.
 2024b) competitions focused on value detection in argumentative texts and multi-lingual traditional media contexts respectively.
-These competitions demonstrated the effectiveness of BERT-based architectures (Schroter, Dementieva, and Groh 2023; Legkas et al.
+CLEF-2024 was the first to identify both endorsed and confronted values in texts.
+These competitions demonstrated the effectiveness of BERT-based architectures (Schroter et al.
+2023; Legkas et al.
 2024) in capturing nuanced value expressions across two different text genres and languages.
-It also demonstrated that the algorithm developed in one context was not successful in identification of values in a second context (Kiesel et al.
+It also demonstrated that algorithms developed in one context may not succeed in identification of values in a another context (Kiesel et al.
 2024b), asserting the importance of contextual models in value detection.
-Generative AI methods were hardly utilized for value detection, with the exception of the FULCRA project (Yao et al.
+Generative AI methods have not yet been utilized for value detection, with the exception of the FULCRA project (Yao et al.
 2024), leveraging LLM-generated text pairs for safety-oriented value detection.
@@
 This novel application to multimodal social media content, combined with our comparative analysis, can contribute valuable insights to both social media analysis and automated value detection research, particularly for platforms where meaning is conveyed through multiple channels simultaneously.
-Figure 1: Values per movie in the annotated data.
+Figure 1: Unique values per TikTok
 Figure 2: Occurrences of each value in the annotated dataset.
-Blue bars indicate the appearance of a value (e.g., achievement) and red bars indicate the conflicted value (e.g., apathy and lack of ambition, conflicting with achievement).
+Blue bars indicate the appearance of an endorsed value (e.g., achievement) and red bars indicate the confronted value (e.g., apathy and lack of ambition, conflicting with achievement).
 3 Data and Annotation
 3.1 Data Curation
-We identified potential TikTok influencers using surveys of children and adolescents (anonymous survey, approved by a the local IRB), and lists provided by three large language models (ChatGPT, Gemini, Bing).
-Two research assistants reviewed the potential influencers, to confirm they are targeting children or adolescents and to classify them by genre.
-This process resulted in a list of 89 influencers, producing content in the major genres prevalent in child- and adolescent-targeted TikTok: beauty and skincare, lifestyle, entertainment and pranks, crafts and DYI, vlogging, gaming and dancing, singing and lip-syncing.
-We sampled 10 TikToks from the page of each influencer.
-We first identified the pinned TikToks, and supplemented the sample with the last TikToks available on March 19, 2024, until a total of 10 TikToks was reached from each influencer, to the total N = 890.
-We only selected TikToks that include verbal sound, such as talking, narration or song. The crawling and curation were done with slight modifications to the API shared by (Teather 2024).
-3.2 Annotation Protocol
-We constructed a codebook on the basis of the Schwartz theory of personal values (Schwartz 1992, 2012).
-The codebook included a definition to each of the 19 values derived from the theory of personal values (Schwartz 2012).
-In addition, the theory included examples of expression of these values that are common on TikTok.
-The codebook included instructions for identifying values as present (1, marked as +), when it is expressed or pursued in the TikTok; as absent (0), when the TikTok does not refer to the value; or as conflicted (-1, marked as -), when the TikTok contradicts or hinders the pursuit of the value.
-Each TikTok could be annotated as referring to as many or as few values as relevant.
-A team of annotators was trained in value annotation.
-Each TikTok was independently annotated by two annotators.
+We identified relevant TikTok influencers using anonymous surveys, asking children and adolescents (local IRB approved) to list the TikTok content creators they follow.
+We also prompted three language models (ChatGPT, Gemini, Bing) to provide lists of prominent influencers with clout over young users.
+Two RAs verified the accounts and confirmed they indeed target children or adolescentsand classified them by genre.
+This process resulted in a list of 91 influencers, active in the following genres: beauty and skincare, lifestyle, entertainment and pranks, crafts and DYI, vlogging, gaming and dancing, singing and lip-syncing.
+We sampled ten TikToks from the page of each influencer, including all TikToks pinned by the influencers, if existed.
+The full dataset contains $N=885$ movies.
+The crawling and curation were done with slight modifications to the Teather API (Teather 2024), collecting only publicly available data.
+Such data curation, handling, processing and analysis, requires careful consideration of ethical aspects and potential harms, from privacy issues to intellectual property rights.
+We discuss potential harms, ethical considerations and handling protocols in Section 7.
+3.2 Data Annotation
+Annotation guidelines were based on the Schwartz theory of personal values (Schwartz 1992, 2012).
+The guidelines provide a definition for each of the 19 values, along with examples of how values are expressed.
+For each TikTok and value, annotators were asked to mark whether the values is endorsed (expressed or pursued in the TikTok; 1), absent (not referenced explicitly or implicitly; 0) or confronted (contradicted or hindered; -1).
+Each TikTok was independently annotated by two annotators from a team trained by the authors.
 A third annotator consolidated the annotations by resolving discrepancies.
-To assess inter-rater agreement on our nominally scaled data, we considered various metrics beyond traditional $\pi$ and $\kappa$ statistics, which are known to be sensitive to imbalanced datasets (Gwet 2008).
-Given that our dataset exhibits significant imbalance, with the majority category (not present) accounting for 0.76 of annotations and rare categories (conflict and dominant) comprising only 0.06 and 0.03, we calculated Gwet’s AC1 coefficient, which has been demonstrated to be more robust against prevalence effects and marginal probability issues than Cohen’s Kappa (Wongpakaran et al.
+The annotation guidelines are available in Appendix A.1.
+We converted each TikTok annotation to a vector of 38 binary labels representing the 19 endorsed values and 19 confronted values.
+The annotated dataset of 885 TikToks contains 2933 (endorsed/confronted) values.
+The distribution of values per TikTok is presented in Figure 1 and the frequency of each value in the dataset is presented in Figure 2.
+The annotated dataset (TikTok IDs and annotations) is available at [URL].
+We used Gwet’s AC1 coefficient to assess inter-rater agreement.
+This coefficient estimates reliability beyond 2 rateres, and has been demonstrated robust against prevalence imbalance and marginal probability issues (Wongpakaran et al.
 2013).
-Our analysis yielded an AC1 score of 0.77, indicating substantial inter-rater agreement despite the challenging data distribution.
-The codebook is available at Appendix A.1.
-In total, the annotated dataset of 890 TikTok TikToks, were annotated with a total of 2968 labels (of values and conflicting values).
-The distribution of values per movie is presented in Figure 1 and the frequency of each value in our dataset is presented in Figure 2.
-We make the annotated dataset available at [URL to added in camera ready version].
+The AC1 agreement was 0.77, indicating substantial inter-annotator agreement.
 4 Computational Approach
-Our computational approach could be placed on three axes: (i) Masked vs.
-Auto-regressive Language Models, (ii) Direct vs.
-indirect, 2-step extraction from TikTok videos (marked d versus s, repestecively), and (iii) Using pretrained LLMs vs.
-fine-tuned models (marked FT).
-In the remainder of this section we describe each axis and the way we integrate them.
-Figure 3 provides a high-level illustration of the architectures.
+We explore a number of computational approaches, differing in two fundamental ways: (i) A direct approach vs.
+a two-step approach, and (ii) Supervised classification vs.
+unsupervised classification using a LLM.
+Direct vs. 2-step Extraction
+We use an LLM as a classifier in two settings: direct and indirect (2-steps).
+In the direct setting, we extract the values directly from the TikTok mp4 video file, either using an LLM, prompting it with the task definition in a few-shots manner, or in a supervised manner, after fusing the video, audio and textual representations (MLMs vs.
+LLMs and Multimodality with Masked Video Autoencoder below).
+The two-step approach allows us to divide and conquer: we first convert a TikTok clip to a full textual script containing master scene heading, action, and dialogue, then extract the values from the textual script.
+The textual scripts are generated by a dedicated prompt.
+This prompt was decided after iterative prompt engineering on a representative subsample of 18 videos.
+Our prompt design prioritized capturing three key video elements: visual content, verbal narration, and text captions.
+Two of the authors independently evaluated script quality, ensuring the conversion accurately captures all three elements.
+The conversion was performed with Gemini 1.5 Pro.
+An example of a generated script is presented in Table 2.
+The prompt we use to convert clips to scripts is available in Appendix A.2.
+Using the generated script, the extraction on the second 2-step could be done either in a supervised manner, including continuous pretraining and fine-tuning, or in a few-shot manner, using the exact same prompt used in the direct approach, albeit, replacing the word ‘movie’ with ‘script’.
+The two pipelines are illustrated in Figure 3.
 Figure 3: Direct and indirect extraction of values from TikTok movies.
-Fine tuning the textual models on scripts is optional.
+continuous pre-training the textual models on scripts is optional.
 LLMs are used in a zero/few-shot manner.
 The MLM must be trained on a subsample in a supervised manner.
-Masked vs. Auto-regressive Language Models
-The classic machine learning framework depends on the availability of a labeled dataset, facilitating training in a supervised manner.
-A Masked Language Model (MLM) is pre-trained to predict words that are masked at random.
-The MLM is then used to create a contextual embedding of an input text, pushing it forward into a classification head.
-In this work we experiment with both roBERTa-large (Liu 2019) and DeBERTaV2-xlarge (He, Gao, and Chen 2023) with an additional classification header performing transfer learning (Devlin 2018).
-We use this supervised approach on textual data only – the scripts and director notes extracted from the videos by another model.
-Generative (auto-regressive) Large Language Models (LLMs) have been demonstrated to perform well on a range of language processing and multimodal tasks.
-LLMs do not require specific training (transfer learning) as classification is done by prompting the model with the task description (Radford et al.
-2019; Brown et al.
-2020; Kojima et al.
-2022).
-State-of-the-Art models such as ChatGPT.4o (OpenAI 2024) and Gemini 1.5 Pro (Google 2024) are multimodal - allowing the user to process different modalities such as text, sound and video.
-In this work we use Gemini 1.5 Pro with a prompt tailored for the specific tasks of script extraction and for direct and 2-step classification (see below).
-In the 2-step setting we also use Mistral 7B (Jiang et al.
+MLMs vs. LLMs
+State-of-the-Art models such as ChatGPT-4o (OpenAI 2024) and Gemini 1.5 Pro (Google 2024) are multimodal, processing different modalities such as text, sound and video.
+Using them allows for a direct approach, analyzing all aspects of the TikTok simultaneously.
+Moreover, the LLMs allow for a few-shot extraction, with no need for supervised training.
+Using an LLM, either in the direct or the 2-step pipeline, the extraction is done by prompting the model with the annotation guidelines, alongside a request to identify the values in a given TikTok (either directly from the mp4 file or from the generated script).
+All prompts are available in Appendix A.
+The LLMs used were Gemini 1.5 Pro and Mistral 7B (Jiang et al.
 2023).
-Specifically, the prompt contains a short formal description of Schwartz’s theory of basic values, a definition for each value, along with some examples and a request to identify the values in a given TikTok (an .mp4 file is uploaded).
-The exact prompts are available in Appendix A.
-Direct vs. 2-step Extraction
-We use an LLM as a classifier in two settings – direct and indirect (2-steps).
-In the direct setting we prompt the model to identify and extract the values directly from the TikTok video.
-In the indirect approach we first prompt an LLM to extract the full script and director’s notes from a TikTok and then extract the values from the script, either in a supervised manner using one of the MLMs or in a few-shot manner, using the exact same prompt we used in the direct approach, albeit, replacing the word ‘movie’ by the word ‘script’.
-An example of a generated script is presented in Table 2.
-We experiment with the different approaches for three main reasons: First, previous work reported results over text in a supervised setting.
-Secondly, the inner-workings of (multimodal) LLMs are opaque and different prompts and tasks may result in unstable results.
-Finally, while LLMs are flashy and do not require supervision, it was argued that they underperform on nuanced tasks such as the detection of emotions, figurative language, implicit hate etc.
-(Ziems et al.
+On the other hand, the use of masked models, either MLMs like BERT or video autoencoders, allows model training, given that annotated data is available.
+The supervised approach is argued to better suit tasks that require nuanced understanding of the content (Ziems et al.
 2024).
+Masked models can be used in both the direct and the two-step pipelines.
+In this work we experiment with roBERTa-large (Liu 2019), DeBERTa-V2-xlarge (He et al.
+2023), and the Longformer (Beltagy et al.
+2020).
+Masked Language Models
+We use pre-trained transformer models with a binary classification head.
+We fine-tuned the classification head with 3-fold cross-validation with early stopping to prevent overfitting.
+We selected the best-performing epoch for each fold based on the F1 score, and then averaged these optimal results across all folds to obtain final performance metrics (commonly referred to as ”best-epoch averaging”, recommended by (Reimers and Gurevych 2017).)
 Genre: Sports/Challenge
@@
 TAYLOR FRITZ is shown on a tennis court serving, then staring into the camera.
-[#9 IN THE WORLD]
-[with one]
-[fastest]
-[serves]
-[240KM/H]
-[I]
-[never]
-[return]
+NARRATOR: #9 IN THE WORLD with one fastest serves 240KM/H I never return
 The NARRATOR is shown holding a tennis racket, about to return a serve.
@@
 NARRATOR: "Oh, come on."
-Table 2: The script generated by the video-to-script prompt (Appendix A.2) for video ID 7341895431883967746
-Transfer learning and Fine-tuning
-Pretrained Language Models, either masked or auto-regressive, serve as foundational models and can be used as-is, either by prompting (LLMs) or by performing transfer learning on a small annotated sample (MLMs).
-However, both types of language models could be further fine-tuned on in-domain data, before the classification phase.
-Typically [^2] Fine-tuning can also be done as part of the classifier training., fine-tuning is done in a similar way to the pretraining, that is – given a dedicated corpora, masked language models are fine-tuned by predicting randomly masked words (Zaken, Goldberg, and Ravfogel 2022), and autoregressive models are fine-tuned on next-word prediction task (Hu et al.
-2021).
-In this work we use the roberta-large framework to fine-tune the roBERTa-large, the DeBERTaV2-xlarge, and the Mistral LLM.
-We report results obtained both by the basic pretrained models and the fine-tuned models.
+Table 2: The script of video ID 7341895431883967746
+Multimodality with Masked Video Autoencoder
+VideoMAE (Tong et al.
+2022), a masked autoencoder pretrained specifically on videos, is reported to achieve impressive results even over small datasets.
+We extract the visual embeddings using VideoMAE, audial embeddings using Wav2Vec2-Base (Baevski et al.
+2020) and the textual embeddings using the SentenceTransformer (Reimers and Gurevych 2019).
+Fusion is done by concatenation.
+These fused embeddings were then fed into an MLP classification head.
+We refer to this fused multimodel F-VideoMAE.
+For more details see Appendix C.2.
+Evaluation Procedure and Metrics
+We evaluate the different pipeline and settings using an the standard F1-score per value and over all values (macro and weighted).
+We also use the recall value in a BLEU-like manner.
+While the different variations of the F-score allows us to evaluate the performance of each system in the detection of each specific value, as well as the overall performance.
+This Recall/BLEU value, inspired by the BLEU (Papineni et al.
+2002) allows us to assess the performance on the movie level, measuring how many of the values conveyed in a specific TikTok were actually recovered by the algorithm, regardless of the specific mixture of values.
+We report and discuss the average Recall/BLEU values.
+In the unsupervised settings (direct and the 2-stage approaches) the results are computed over the whole dataset.
+Settings involving supervised training were conducted in a 3-fold cross-validation manner, each fold uses two thirds of the data for training and a third for testing.
+Results are reported only for values appearing in at least 5% of the movies, in order to allow enough data in the train/test split.
 Compute Resources
-All experiments and training were executed on Google Colab Enterprise runtime environment with an NVIDIA Tesla A100 GPU (a2-highgpu-1g configuration).
-5 Results, Analysis, and Discussion
-The F-scores achieved by each of the models on each of the nineteen values are presented in Figure 4 and the weighted and macro F-scores are presented in Table 3.
-The best overall results for values and their conflicted counterparts were achieved with the non-fine-tuned RoBERTa-large (roBERTas) with weighted F-scores of 0.49/0.42/0.47 (present/contradicted/all [$+/-/all$], respectively).
-These levels of classification reflect the opacity of the signal, that corresponds to a multidimensional and complex human phenomenon.
-In fact, even human annotators required extensive training and still found the task non-trivial.
-At the same time, these results are competitive with previously reported state-of-the-art results (Schroter, Dementieva, and Groh 2023; Legkas et al.
-2024) achieved in less challenging settings in which (i) text, rather than video, was the original medium of the input, and (ii) the number of examples available for training was an order of magnitude larger.
-[^3] A direct comparison is unattainable due to the very different nature of the data in terms of source, medium and style.
-Looking at specific values (Figure 4(a)), best classification was achieved for Achievement (0.71), Benevolence Dependability (0.61), and Security Personal (0.58), all by the RoBERTas model.
-Achievement, present in more than one third of the videos, is the most frequent value in the annotated dataset, thus more occurrences are found in training and in the test set.
-However, the two other values occur far less than most values, hence performance cannot be attributed to prior frequency alone.
-A fine-grained analysis of the results by method and value reveals interesting patterns, suggesting a number of insights. We address some of these patterns in the remainder of this section.
-(a) Present manifestations (Max. F-Score: 0.71)
-(b) Contradicting manifestations (Max. F-Score: 0.54)
-Figure 4: F-Scores per each model and value. Values appearing less than 30 times were not included in the BERT based MLM models, thus excluded from the analysis (appearing with score 0 in the figure).
-Masked Language Models Large Language Models
+All experiments were executed on Google Colab Enterprise runtime environment with an NVIDIA Tesla A100 GPU (a2-highgpu-1g configuration).
+Endorsed Confronted All
+W M W M W M
+F-VideoMAEd 0.23 0.20 0.22 0.17 0.23 0.19
+Geminid 0.46 0.40 0.26 0.27 0.4 0.37
 […]
-Weighted+ 0.49 0.44 0.30 0.33 0.29 0.33 0.40
-Macro+ 0.44 0.38 0.26 0.29 0.33 0.28 0.33
-Weighted- 0.42 0.41 0.33 0.30 0.17 0.21 0.22
-Macro- 0.35 0.35 0.27 0.25 0.22 0.18 0.22
-Weightedall 0.47 0.43 0.31 0.33 0.26 0.30 0.35
-Macroall 0.39 0.36 0.25 0.26 0.29 0.24 0.29
-Table 3: Weighted and Macro F-Score results for the values ($+$), the conflicted values (--) and all ($all$).
-$FT$: fined-tuned model; $s$ indicates extraction from scripts (2-step extraction); $d$ indicates direct extraction from video.
-MLMs vs. LLMs
-The best LLM results were obtained by Geminid.
-However, it was outperformed by both RoBERTas and DeBERTas in all metrics and over the three partitions ($\{weighted,macro\}\times\{+,-,all\}$, Table 3).
-Looking at single values, we observe that Gemini outperforms the MLMs in cases in which the number of occurrences of the value is minimal, hence supervised training is hampered, see results per value in Figure 4 and compare to the value distribution Figure 2.
-An example is Universalism Concern in the present manifestations, and Power Resources in the contradiction manifestation.
-These results are consistent with the literature, reporting that strong supervised models tend to outperform current LLMs on tasks requiring nuance, typical to the Computational Social Science (CSS) research (Ziems et al.
-2024).
-Pretraining vs. Fine-tuning
-All the fine-tuned models fell short, comparing to their pretrained counterparts.
-While somewhat surprising, previous studies report similar trends in various tasks and settings, e.g., (Zaken, Goldberg, and Ravfogel 2022; Kumar et al.
-2022), as well as in the best performer in she value detection shared task (Schroter, Dementieva, and Groh 2023).
-Given the task and data at hand, we attribute this result to the combination of two main factors: (i) The relatively small dataset used for finetuning, (ii) The large number of imbalanced classes, as presented in Figure 2.
-Future work will explore the impact of each of these factors.
-Direct vs. 2-step Extraction
-The textual models (MLMs) applied in a 2-step pipeline and trained in a supervised manner, achieved better results than Geminid applied directly on the videos.
-However, comparing apples to apples, considering only the LLMs, we found that the direct application of Geminid on the videos significantly outperformed the application of both Geminis and Mistrals on the scripts in the 2-step pipeline.
-These results highlight the importance of the visual and sound elements to the TikTok platform, and thus the potential of the direct multimodal approach.
-Present vs. Conflicting Manifestations
-Most present values are more frequent (Figure 2) and better detected than values in conflict (Figure 4 and Table 3).
-It is interesting to note that while Geminid, our top performing LLM, showed decent performance on the present values (Weighted+ F-score of 0.4, compared to 0.49 by the top performing model and 0.33 by the second best LLM), it perform poorly on the conflicting values, achieving an F-score of only 0.22, compared to the 0.42 by the top performing model and only slightly better than the other LLMs.
-This result is surprising: while MLM results are expected to drop due to the smaller number of labeled instances available for training, we would have expected LLMs, used in a few-shot manner, to perform comparably on present and conflicting values.
-We have no plausible explanation for this significant drop in performance and intend to further explore it in the future.
-Video to Script Conversion
-Finally, the use of the multimodal Gemini 1.5Pro to convert videos to detailed textual script (master scene heading, action, dialogue) proved beneficial for the task.
-We find the conversion to be reasonably accurate and the level of details provide useful markers.
-This conversion was vital for the application of the successful MLM.
-Two points are worth highlighting: First, before converging on the final prompt (see Appendix A.2) we have experimented with different prompts, producing different outputs in various formats and detail.
-The evaluation of the quality of the script was done in a qualitative manner by two of the authors.
-Second, even the state-of-the-art models often fail to capture important details.
-For example, in a humerus nonsensical TikTok showing the influencer air-driving a car as a part of a guessing charade game, practically air driving, floating around in the room and bumping into objects, the generated script failed to reflect the fact that the influencer was actually moving in the air – the main point of the video.
-Limitations
+[…]
+Longformer2s 0.34 0.30 0.28 0.20 0.33 0.27
+DeBERTa2s 0.36 0.33 0.35 0.32 0.36 0.33
+Mistral2s 0.39 0.36 0.26 0.26 0.36 0.33
+RoBERTa2s 0.42 0.38 0.38 0.35 0.41 0.37
+Gemini2s 0.48 0.43 0.36 0.37 0.45 0.41
+Table 3: Weighted (W) and Macro (M) F-Score results for the Endorsed Values , the Confronted Values and All.
+The direct and two-steps settings are indicated by $d$ and $2s$ (for 2-step), respectively; PTPT indicates continuous Pre-Training.
+(a) Endorsed Values
+(b) Confronted Values
+Figure 4: F-Scores per each model and value.
+Values appearing less than 5% of TikToks were not included in the BERT based MLM models, thus excluded from the analysis (appearing with score 0 in the figure).
+5 Results and Analysis
+Overall Results
+Table 3 presents the weighted and the macro F-scores achieved by each of the models.
+Best results overall were obtained by Gemini2s (using the 2-stage pipeline, prompting Gemini in both stages) with 0.45 (0.41) weighted (macro) F-score.
+Separating endorsed and confronted values, this setting achieves best results (weighted and macro) on the endorsed values, and best macro results on the confronted values.
+While the RoBERTa2s achieved the highest weighted F-score on the confronted values, Gemini2s is a close second, with 0.36, compared to 0.38 by RoBERTa2s.
+Interestingly, the direct pipeline (Geminid) tend to underperform, compared to the use of the Gemini or the RoBERTa in a 2-stage pipeline.
+These results suggest that while Gemini was trained in a multimodal manner over text and video, nuanced classification better done over texts.
+Indeed, the 2-stage approach harnesses the multimodal capabilities of Gemini to convert videos to detailed textual scripts (master scene heading, action, dialogue) and then harness the capabilities of the LLM as a strong few-shot learner.
+The detailed script generated in the first stage helps mitigating the challenges demonstrated by Ziems et al.
+(2024).
+Results per Movie (Recall/BLEU)
+Focusing on the movie level, rather than on the value level, the 2-stage pipeline proved to achieve the best results (Table 4).
+This is consistent with the results presented above.
+Gemini2s demonstrates superior performance on the endorsed values and achieves the best overall results.
+However, when focusing specifically on confronted values at the movie level, RoBERTa2s outperforms Gemini2s.
+We attribute this gap to the fact that confronted values are more opaque and therefore harder to detect without explicit supervision.
+At the movie level, evaluation is based on recall rather than F1; for confronted values, which are more opaque, this favors models such as RoBERTa2s that exhibit higher sensitivity.
+This increased sensitivity, however, is accompanied by a substantially higher rate of false positives, which leads to lower precision, as evident at the value level (Table 6).
+However, at the movie level, where evaluation emphasizes recall rather than precision, this reduction in precision is not directly reflected.
+In summary, for the confronted values, at the value level this results in comparable F1 scores to Gemini2s (Table 3), while at the movie level, the higher recall of RoBERTa2s leads to improved performance.
+Model Endorsed Confronted All Values
+Geminid 0.51 0.2 0.36
+Mistral2s 0.51 0.26 0.39
+[…]
+[…]
+Gemini2s 0.58 0.32 0.45
+DeBERTa2s 0.39 0.39 0.39
+roBERTa2s 0.43 0.44 0.43
+Table 4: Recall/BLEU results. The direct and two-steps pipelines are indicated by $d$ and $2s$ (2-step), respectively; PTPT indicates continuous Pre-Training.
+Results per Value
+The results per value are presented in the radar plots in Figure 4.
+The highest F-score for endorsed values (0.69) was obtained for Achievement, followed by Security Personal and Power Resources (0.54 and 0.53, respectively) while the highest score for a confronted value (0.48) was obtained for Humility, followed by Security Personal and Face (0.45, 0.41).
+Indeed, these are three of the five most frequent confronted values, see Figure 2.222We note that the numbers reported here result from different settings, see detailed results in Appendix B.
+Endorsed vs. Confronted Values
+Most of the endorsed values are more frequent in the dataset (Figure 2) and were better detected than confronted values (Figure 4 and Table 3).
+All models showed a performance drop between the endorsed and confronted values.
+The three top performing models (pipelines), Geminid, RoBERTa2S, and Gemini2s, exhibited an interesting pattern.
+The performance drop for Geminid (direct pipeline) was nearly twice as large as that of Gemini2s, which itself was about three times the drop observed for RoBERTa2s.
+The weighted F1-scores over the endorsed values were 0.48, 0.46, and 0.42 for Gemini2s, Geminid, and RoBERTa2s, respectively.
+Over the confronted values, the F1-scores decreased to 0.36 ($\Delta=0.12$), 0.26 ($\Delta=0.20$), and 0.38 ($\Delta=0.04$), respectively.
+This drop in performance for confronted values is also evident at the movie level, where we aggregate over all values per video (Table 4).
+Gemini2s maintains the best overall performance (0.45), while RoBERTa2s outperforms Gemini2s on the confronted values (0.44 vs.
+0.32), reflecting its higher sensitivity in detecting these opaque manifestations.
+The direct pipeline Geminid suffers a large drop (0.20), consistent with the pattern observed at the value level (Table 3).
+It seems the manifestations of confronted values pose a challenge even for trained annotators: their detection requires an understanding of the meaning of the value, the meaning of its absence, and the way in which its absence (or confrontation) is promoted in an ironic or subversive manner.
+These results suggest that LLMs fall short in processing conceptual complexity and social opposition.
+However, these observations highlight the advantage of the 2-stage pipeline: by combining multimodal LLMs in a divide-and-conquer manner, it mitigates some of their shortcomings, achieving more robust performance both at the value level and when aggregating per movie.
+Continuous Pretraining
+Continuous pretraining did not improve any of the models we used.
+We attribute this to the combination of two main factors: (i) The relatively small dataset used for the continuous pretraining (<<1000 videos), (ii) The large number of unbalanced classes, as presented in Figure 2.
+Future work should explore the impact of each of these factors.
+Limitations and Discussion
 This study is not without limitations.
-First, we sampled influencers in seven specific genres.
-In our dataset, we found much variability between the genres in content as well as style.
-We identified genres most important for children and adolescents, yet know that TikTok is extremely varied, and other genres exist.
-Thus, it is yet to verify that our results generalize across TikTok.
-Second, we applied the Schwartz personal values theory to define and identify values.
+First, we have sampled 91 influencers and 885 videos from TikTok.
+Our manual annotation process, included two independent expert annotators, and a third consolidator.
+This process ensured quality, yet limited the size of the dataset and the generalization beyond the TikTok platform.
+While this work explores the different pipelines and methods for value extraction, future work will incorporate a larger dataset, allowing for better generalization even for values infrequent in the current dataset.
+Second, we have sampled influencers from seven specific genres.
+In our dataset, we found much variability across the genres in content and style.
+The genres we explored were identified as the most important for children and adolescents.
+TikTok, however, is extremely diverse and other genres may become prominent.
+Thus, future studies must verify that our results generalize across TikTok.
+A third issue stems from the movie-to-script conversion.
+The quality of the generated scripts was established in a qualitative manner by two of the authors.
+However, even state-of-the-art models may fail to capture important details.
+For example, a humerus nonsensical TikTok showed the influencer air-driving a car in a charade game, appearing to be practically driving in the air, floating around the room and bumping into objects.
+The generated script failed to reflect the fact that the influencer was depicted as actually moving in the air.
+Forth, we applied the Schwartz personal values theory to define and identify values.
 This theory was tested across cultures and in hundreds of samples across the world (Sagiv and Schwartz 2022), and is considered the leading theoretical framework of values in social science.
@@
 2025).
-Last, in our analysis, we utilized two MLMs, and two LLMs.
-the use of other models may offer different results, and future studies can extend this investigation further.
+Finally, while this work explored the use of an array of strong models – new models, announced periodically, may offer a different perspective (e.g., significant improvement of the direct pipeline).
 6 Conclusions and Implications
-We present pipeline for the extraction of personal values from TikTok, and demonstrate their utility.
-We show that (supervised) BERT based models outperform LLMs in value extraction, and offer a method for conversion of multimodal TikToks into text appropriate for BERT training and analysis.
-We also show variability in detection between different values, and between values supported and those contradicted in the TikToks.
+We have explored two pipelines for the extraction of personal values from TikTok – a direct approach using a SOTA multimodal LLMs and an indirect, two-steps, approach using either multimodal LLM for both stages or an LLM in the first step and a supervised MLM in the second step.
+We have demonstrated that the two-steps approach mitigates some of the shortcomings of SOTA LLMs – particularly when it comes to the processing of nuanced societal commentary and values.
+We further analyze the performance over individual values and over endorsed and confronted values.
+The results reported in this paper outperform state-of-the-art results (Schroter et al.
+2023; Legkas et al.
+2024) achieved in less challenging settings in which (i) text, rather than video, was the original medium of the input, and (ii) the number of examples available for training was an order of magnitude larger.
+[^3] A direct comparison is unattainable due to the very different nature of the data in terms of source, medium and style.
 The methods presented have important implications for research and society.
-TikTok emerged as a leading platform among youth (Anderson, Faverio, and Gottfried 2023; Woodward 2024), yet at the same time produces worry among parents, educators, and decision makers (Conte et al.
-2024).
-By applying value detection, research can move on to investigate value transmission in TikTok at scale.
+TikTok emerged as a leading platform among youth (Anderson et al.
+2023; Woodward 2024), yet at the same time produces worry among parents, educators, and decision makers (Conte et al.
+2025).
+By applying value detection, researchers can study the prevalence transition, and effect of different values in TikTok, at scale.
 Such research can provide society with much needed information regarding the social environment of young people.
+Importantly, rather than approaching social media influence as a monolithic concern, parents can use our value identification framework to have more nuanced conversations with their children about the content they consume.
+In addition, the methods presented here have important implications for AI alignment, platform governance, and design choices.
 It is important to note that advertisers and influencers can make similar use of the method, in order to target youth more effectively.
-We follow (De Veirman, Hudders, and Nelson 2019) in call for policy makers and educators to mitigate misuse of online platforms by enabling youth to better identify patterns of influence in social media.
+We follow De Veirman et al.
+(2019) in call for policy makers and educators to mitigate misuse of online platforms by enabling youth to better identify patterns of influence in social media.
+7 Ethical Considerations
+Social media platforms create a semi-public sphere where influencers actively post content not just for their friends, but for a wider audience.
+Influencers often use their online presence as their main source of income, putting vast effort to maximize exposure and engagement (Carpiano et al.
+2023; DiResta 2024).
+Therefore, studying influencer culture and ecosystem is not only ‘fair-practice’ under the unobtrusive observer paradigm (Kozinets 2010), but also crucial for understanding how values are being monetized by influencers targeting children.
+We note that eight accounts in our data feature minors, three of these are parents accounts.
+The legal guardians are active participants and supporters of the other five accounts.
+The dataset curated for this work is publicly available and collected only via public APIs with no violation of TikTok terms of service.
+The annotated dataset available at [URL] contains only the TikTok IDs and the annotations, in order to allow influencers maintaining privacy should they choose to delete their content.
+The societal need to understand the ecosystem that targets and influences minors balances and mitigates these ethical concerns.
+Moreover, we believe that only such understanding will allow parents, educators, and policy makers, to support minors – a vulnerable population – in facing potential risks resulting from the broad use of social media.
 Appendix A Prompts and Responses
@@
 This very same “prompt” served as the annotation guidelines (code book) offered to annotators.
-Note that the figure contains an examples (“few shot”) for only one value (SELF-DIRECTION-THOUGHT) while the actual prompt (and the annotation guidelines) provided examples for each of the values.
+Note that the figure contains an example (“few shot”) for only one value Self-direction Thought while the actual prompt (and the annotation guidelines) provided examples for each of the values.
 The same prompt was used for few-shot extraction from scripts in the 2-step extraction, the sole difference is that the opening line is “You are given a movie script of a TikTok video”, rather than “You are given a TikTok video”.
-The full code book (prompt) is available at [add URL].
+The full code book (prompt) will be available at [URL to be added in camera ready version].
 Figure 5: Annotation guidelines and prompt for direct extraction.
 Figure 6: Additional guidelines, appended after all the examples (code book).
-A.2 Prompt for Srtipt Extraction
+A.2 Prompt for Script Extraction
 The prompt used to convert TikTok movies to textual scripts is available in Figure 7; the script generated by applying this prompt on TikTok ID 7341895431883967746 is presented in Table 2.
-The manual annotations for this specific movie are: Achievement- present (1), Face - conflict (-1).
+The manual annotations for this specific movie are: Achievement Endorsed Value (1), Face Confront value (-1).
 Figure 7: The prompt used to convert TikTok movies to textual scripts.
+Appendix B Detailed Results
+Detailed results per value are provided in Tables 5 and 6.
+Appendix C Experimental Design: configurations and hyper-parameters
+In this section we define the hyper-parameters used in this paper.
+C.1 MLM Training Configuration
+All MLMs were trained using the AdamW optimizer with a learning rate of 2e-5, following standard practices for transformer fine-tuning.
+We employed a linear learning rate schedule with warm-up, where the number of warm-up steps was set to 10% of the total training steps, allowing the model to gradually adapt to the task-specific data distribution and preventing early training instability.
+We used 5 epochs with early stopping with patience of 2 epochs.
+Throughout training, transformer backbone weights remained unfrozen, enabling full model adaptation to the nuances of human values classification rather than relying solely on pre-trained representations.
+C.2 Multimodal Extraction using VideoMAE
+Visual embeddings were extracted by sampling 16 uniformly distributed frames from each video using PyAV and processed through VideoMAE-Base to obtain visual representations.
+Audio embeddings were generated by extracting 10-second audio segments using librosa (McFee et al.
+2015) and processing through Wav2Vec2-Base (facebook/wav2vec2-base) with mean pooling over the temporal dimension to produce 768-dimensional feature vectors.
+Text embeddings were obtained by performing Optical Character Recognition (OCR) using Tesseract on three key frames (first, middle, and last) to extract visual text overlays, concatenating the extracted text, and encoding it using the SentenceTransformer model ’all-MiniLM-L6-v2’ to generate 384-dimensional semantic representations.
+The embeddings of the three modalities were then fused by concatenation to form comprehensive multimodal representations with a total dimensionality of 768 (VideoMAE) + 768 (Wav2Vec2) + 384 (SentenceTransformer) = 1,920 features per video.
+[…]
+Value P R F1 P R F1 P R F1 P R F1 P R F1 P R F1 P R F1 P R F1 P R F1
+Self-Direction Thought 0.49 0.33 0.40 0.47 0.36 0.40 0.32 0.61 0.42 0.43 0.46 0.44 0.45 0.47 0.46 0.28 0.35 0.31 0.39 0.33 0.36 0.23 1.00 0.37 0.12 0.15 0.13
+Self-Direction Action 0.43 0.67 0.53 0.49 0.56 0.52 0.34 0.75 0.47 0.49 0.45 0.47 0.37 0.41 0.38 0.27 0.30 0.28 0.40 0.44 0.41 0.26 1.00 0.41 0.31 0.21 0.22
+Stimulation 0.34 0.83 0.48 0.36 0.78 0.49 0.33 0.70 0.45 0.42 0.36 0.38 0.34 0.33 0.34 0.29 0.35 0.31 0.33 0.35 0.34 0.25 1.00 0.40 0.10 0.19 0.13
+Hedonism 0.23 0.93 0.37 0.27 0.87 0.41 0.26 0.80 0.39 0.31 0.45 0.37 0.25 0.38 0.30 0.17 0.32 0.22 0.17 0.31 0.22 0.17 1.00 0.29 0.25 0.19 0.21
+Achievement 0.61 0.77 0.68 0.68 0.71 0.69 0.57 0.62 0.59 0.70 0.56 0.62 0.49 0.46 0.47 0.33 0.29 0.31 0.47 0.40 0.43 0.36 1.00 0.53 0.44 0.47 0.46
+Power Dominance 0.46 0.47 0.46 0.51 0.36 0.42 0.39 0.27 0.32 0.34 0.44 0.37 0.25 0.38 0.30 0.25 0.38 0.30 0.24 0.35 0.28 0.19 0.99 0.32 0.23 0.24 0.23
+Power Resources 0.69 0.43 0.53 0.68 0.32 0.44 0.47 0.28 0.35 0.29 0.36 0.32 0.29 0.35 0.31 0.23 0.32 0.26 0.29 0.33 0.31 0.20 1.00 0.33 0.12 0.40 0.17
+Face 0.16 0.52 0.24 0.16 0.25 0.19 0.11 0.22 0.14 0.16 0.48 0.24 0.09 0.25 0.13 0.09 0.27 0.13 0.09 0.21 0.12 0.08 1.00 0.15 0.00 0.00 0.00
+Security Personal 0.35 0.83 0.50 0.39 0.87 0.54 0.35 0.67 0.46 0.53 0.53 0.48 0.54 0.59 0.52 0.11 0.34 0.16 0.10 0.27 0.15 0.10 1.00 0.17 0.15 0.37 0.21
+Security Social 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA
+Tradition 0.35 0.55 0.43 0.29 0.45 0.35 0.20 0.39 0.27 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA
+Conformity Rules 0.15 0.46 0.23 0.12 0.38 0.18 0.11 0.46 0.18 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA
+Conformity Interpersonal 0.40 0.20 0.27 0.40 0.20 0.26 0.10 0.12 0.10 0.19 0.45 0.26 0.15 0.38 0.21 0.13 0.38 0.19 0.10 0.27 0.15 0.10 1.00 0.18 0.10 0.78 0.18
+Humility 0.35 0.14 0.21 0.00 0.00 0.00 0.19 0.14 0.16 0.10 0.43 0.17 0.10 0.46 0.16 0.06 0.32 0.10 0.11 0.49 0.18 0.07 1.00 0.13 0.07 0.09 0.08
+Benevolence Care 0.63 0.486 0.53 0.50 0.42 0.46 0.53 0.37 0.43 0.30 0.39 0.34 0.29 0.36 0.32 0.26 0.33 0.29 0.23 0.28 0.25 0.21 1.00 0.35 0.23 0.48 0.26
+Benevolence Dependability 0.43 0.39 0.41 0.46 0.34 0.38 0.34 0.35 0.35 0.55 0.45 0.48 0.57 0.38 0.42 0.14 0.24 0.17 0.23 0.37 0.28 0.16 1.00 0.28 0.24 0.65 0.32
+Universalism Concern 0.33 1.00 0.50 0.25 1.00 0.40 0.03 1.00 0.05 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA
+Universalism Nature 0.67 0.22 0.33 0.50 0.11 0.18 0.10 0.33 0.16 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA
+Universalism Tolerance 0.33 0.20 0.25 0.20 0.20 0.20 0.05 0.40 0.10 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA
+Table 5: Endorsed Values: Precision, Recall and F-scores for each model and value.
+Values marked with indicate categories with fewer than 5% of instances in the dataset.
+P = Precision, R = Recall, F1 = F1-score PTPT: continuous pre-trained model; $2s$ indicates 2-step extraction (extraction from scripts); $d$ indicates direct extraction from video.
+[…]
+Value P R F1 P R F1 P R F1 P R F1 P R F1 P R F1 P R F1 P R F1 P R F1
+Self-Direction Thought Neg 0.09 0.08 0.09 0.00 0.00 0.00 0.02 0.17 0.04 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA
+Self-Direction Action Neg 0.19 0.25 0.22 0.20 0.12 0.15 0.11 0.29 0.16 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA
+Stimulation Neg 0.00 0.00 0.00 0.00 0.00 0.00 0.01 0.33 0.02 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA
+Hedonism Neg 0.08 0.50 0.13 0.06 0.17 0.09 0.04 0.67 0.07 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA
+Achievement Neg 0.30 0.68 0.42 0.38 0.42 0.40 0.04 0.26 0.07 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA
+Power Dominance Neg 0.30 0.22 0.26 0.31 0.13 0.19 0.05 0.13 0.07 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA
+Power Resources Neg 0.27 0.33 0.30 0.44 0.44 0.44 0.01 0.11 0.01 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA
+Face Neg 0.41 0.40 0.41 0.40 0.22 0.29 0.20 0.34 0.25 0.20 0.50 0.28 0.16 0.4 0.23 0.09 0.29 0.13 0.13 0.33 0.18 0.10 1.00 0.18 0.00 0.00 0.00
+Security Personal Neg 0.39 0.53 0.45 0.40 0.46 0.43 0.22 0.39 0.28 0.18 0.55 0.27 0.20 0.57 0.30 0.08 0.29 0.13 0.10 0.32 0.15 0.10 1.00 0.18 0.09 0.07 0.08
+Security Social Neg 0.20 0.33 0.25 0.50 0.44 0.47 0.03 0.33 0.05 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA
+Tradition Neg 0.00 0.00 0.00 0.50 0.25 0.33 0.00 0.00 0.00 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA
+Conformity Rules Neg 0.07 0.38 0.11 0.10 0.46 0.17 0.04 0.38 0.07 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA
+Conformity Interpersonal Neg 0.43 0.27 0.33 0.56 0.18 0.27 0.32 0.34 0.33 0.30 0.52 0.38 0.27 0.47 0.34 0.16 0.29 0.20 0.16 0.28 0.20 0.13 1.00 0.23 0.18 0.76 0.25
+Humility Neg 0.43 0.32 0.37 0.58 0.15 0.24 0.27 0.20 0.23 0.52 0.45 0.48 0.42 0.41 0.41 0.34 0.36 0.35 0.33 0.34 0.33 0.24 1.00 0.39 0.38 0.44 0.41
+Benevolence Care Neg 0.68 0.17 0.27 0.56 0.08 0.14 0.22 0.21 0.22 0.26 0.56 0.35 0.26 0.46 0.33 0.17 0.37 0.23 0.17 0.34 0.23 0.00 0.00 0.00 0.34 0.07 0.09
+Benevolence Dependability Neg 0.27 0.24 0.26 0.42 0.12 0.19 0.04 0.10 0.06 0.26 0.56 0.35 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+Universalism Concern Neg 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA
+Universalism Nature Neg 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA
+Universalism Tolerance Neg 0.10 1.00 0.18 0.10 1.00 0.18 0.01 1.00 0.02 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA
+Table 6: Confronted Values: Precision, Recall and F-scores for each model and value.
+P = Precision, R = Recall, F1 = F1-score PTPT: continuous pre-trained model; $2s$ indicates 2-step extraction (extraction from scripts); $d$ indicates direct extraction from video.